# GDEPO: Group Dual-dynamic and Equal-right Advantage Policy Optimization with Enhanced Training Data Utilization for Sample-Constrained Reinforcement Learning

### v2

@@ I Introduction
Figure 1: LLMs serve as auto-theorem provers by receiving a standardized prompt containing $G$ identical problem instances, generating whole-proof completions, and verifying them through a Lean4 server [20]. A sample is deemed invalid if all outputs in the group are incorrect. For the same prompt, sampling continues up to a maximum number of attempts until at least one correct proof is obtained. When the within-group accuracy falls below a predefined threshold, the correctness feedback from the verifier contributes a sign term to the advantage function, while other signals determine its magnitude. For samples initially classified as invalid but later yielding valid proofs through subsequent sampling, an additional backward pass is performed to refine the model parameters.
@@ V-B Utilization Effect of Challenging Training Data Evaluation
Figure 4: Dae. Ad quatur autat ut porepel itemoles dolor autem fuga. Bus quia con nessunti as remo di quatus non perum que nimus. (a) Case I. (b) Case II.
@@ V-C Effectiveness of Equal-Right Advantage Evaluation
To assess the contribution of the equal-right advantage module to the effective utilization of composite rewards, we compare three variants: GRPO with a single binary reward, GRPO with composite rewards, and our proposed GDEPO. All three algorithms are trained on the same dataset of 30,000 problems containing 30% challenging samples. The composite reward function used in GRPO matches GDEPO exactly in both reward types and weights to eliminate any confounding effects from reward weighting. As shown in Table II, GDEPO reduces both the average output length and repetition rate without compromising success rate, by properly computing the advantage based on composite rewards. In the table, “number of tokens” denotes the average token count across all completions, and “similarity rate” refers to the average 5-gram repetition rate. While directly applying composite rewards in GRPO also suppresses output length and redundancy, it incurs a noticeable drop in correctness—a critical metric. In contrast, GDEPO preserves correctness by using it to determine the sign of the advantage function, thereby ensuring a proper policy optimization direction, while leveraging auxiliary rewards to modulate the magnitude of policy updates and balance multiple objectives.
@@
The performance difference relative to Goedel-Prover-V2-8B with GRPO (Single Binary Reward) is indicated as $\delta$ Perf.”. Tokens = Number of Tokens, Simi. Rate = Similarity Rate, Prob. Solved = Number of Problem Solved.

### v3

@@ I Introduction
Figure 1: The detailed training procedure of GDEPO: LLMs serve as auto-theorem provers by receiving a standardized prompt containing $G$ identical problem instances, generating whole-proof completions, and verifying them through a Lean4 server[20]. A sample is deemed invalid if all outputs in the group are incorrect. For the same prompt, sampling continues up to a maximum number of attempts until at least one correct proof is obtained. When the within-group accuracy falls below a predefined threshold, the correctness feedback from the verifier contributes a sign term to the advantage function, while other signals determine its magnitude. For samples initially classified as invalid but later yielding valid proofs through subsequent sampling, an additional backward pass is performed to refine the model parameters.
@@ V-B Utilization Effect of Challenging Training Data Evaluation
Figure 4: Performance comparison of the same base model trained with GRPO using a single binary reward versus GDEPO employing composite rewards, under varying proportions of challenging samples: (a) performances on MathOlympiadBench; (b) performances on PutnamBench.
@@ V-C Effectiveness of Equal-Right Advantage Evaluation
To assess the contribution of the equal-right advantage module to the effective utilization of composite rewards, we compare three variants: GRPO with a single binary reward, GRPO with composite rewards, and our proposed GDEPO. To ensure the best performance of the algorithms, GDEPO is trained on 30,000 training samples containing 30% challenging samples, while GRPO with two different reward configurations is trained on 30,000 samples containing 20% challenging samples. The composite reward function used in GRPO matches GDEPO exactly in both reward types and weights to eliminate any confounding effects from reward weighting. As shown in TableII, GDEPO reduces both the average output length and repetition rate without compromising success rate, by properly computing the advantage based on composite rewards. In the table, “number of tokens” denotes the average token count across all completions, and “similarity rate” refers to the average 5-gram repetition rate. While directly applying composite rewards in GRPO also suppresses output length and redundancy, it incurs a noticeable drop in correctness—a critical metric. In contrast, GDEPO preserves correctness by using it to determine the sign of the advantage function, thereby ensuring a proper policy optimization direction, while leveraging auxiliary rewards to modulate the magnitude of policy updates and balance multiple objectives.
@@
The performance difference relative to Goedel-Prover-V2-8B with GRPO (Single Binary Reward) is indicated as $\delta$ Perf.”. Tokens = Number of Tokens, Simi. Rate = Similarity Rate, Prob. Solved = Number of Problem Solved. Higher values are preferred for metrics marked with $\uparrow$, and lower values are preferred for those marked with $\downarrow$.